An outpatient who is pregnant collected a urine sample for creatinine clearance as her physician ordered. But, she collected twenty hours because she thinks it would not matter if the collection time is different. She provided her height and weight (5’7” and 100 kilograms). The lab assistant, who received her urine, measured the total volume and reported as 3,500 ml, made an aliquot then she discarded the rest of the urine. The blood was centrifuged and was sent to chemistry for analysis of plasma creatinine and the urine aliquot for urine creatinine. The blood result was 1.8 mg/dl and the urine creatinine was reported 175 mg/dl. Calculate and report the actual creatinine clearance in ml/min.
Answer #1

Calculation of actual creatinine clearance:

Step 1:

Formula for creatinine clearance

CrCl = UCr * UVol
filler.gif
PCr * Tmin

Step 2:

Ucr Urine creatinine = 175mg/dL

Uvol Urine volume = 3500mL.

Pcr Plasma creatinine = 1.8mg/dl

Tmin Total collection time = 20 hours

To calculate in min = 20 X 60

Step 3:

By applying into the formula,

CrCl = \frac{175\times 3500}{1.8 \times 20\times 60}

= \frac{612500}{2160}

CrCl = 283.6 mL/min.

Step 4:

Corrected CrCl = CrCl \times \frac{1.73}{BSA}

Step 5:

Calculate BSA by using Mosteller formula

BSA = \sqrt{height in cm X weight in kg/3600}

Step 6:

Convert height into centimeter

1 foot = 30.48 cm.

5'7" = 173.7cm.

Weight = 100kg

Step 7:

By applying into the BSA formula

BSA = \sqrt{173.7X 100/ 3600}

BSA = 2.19m2

Step 8:

Calculate the corrected creatinine clearance

CrCl = 283.6 X \frac{1.73}{2.19}

Corrected Creatinine clearance = 224 mL/min.

The actual creatinine clearance is 224mL/min.